OpenStudy (anonymous):

f(x)=(2x+3)^−2, f'(x)=?

hartnn (hartnn):

do you know chain rule ?

OpenStudy (anonymous):

i got -2(2x+3^1*(2)

OpenStudy (anonymous):

-2(2x+3)^-1*2

OpenStudy (anonymous):

yes, I have learnt chain rule just before.

hartnn (hartnn):

how did the exponent become -1 ?? -2-1 =... ?

OpenStudy (anonymous):

almost right, but the exponent rules says that n-1, in this case n=-2, so then -2-1=-3

OpenStudy (anonymous):

AHHH I see!!

hartnn (hartnn):

everything else is correct...

OpenStudy (anonymous):

Thank you. w(r)=sqrt(r^a+1) I have got 1/2(r^a+1)^(-1/2) where did it miastake?

OpenStudy (anonymous):

sorry i mean a is 9

hartnn (hartnn):

chain rule says to diff. inner function also, so here, derivative of r^a +1 will get multiplied.

hartnn (hartnn):

1/ [2 sqrt(r^9+1)] [d/dx (r^9+1)]

hartnn (hartnn):

(1/2) (r^9+1)^(-1/2) * (.....)

OpenStudy (anonymous):

9r?

hartnn (hartnn):

for x^n its n x^{n-1} for r^9, it'll be .... ? here, x=r, n=9

OpenStudy (anonymous):

9r^8?

hartnn (hartnn):

thats correct :)

hartnn (hartnn):

(1/2) (r^9+1)^(-1/2) * (9r^8)

OpenStudy (anonymous):

thank you, one more question how do you do f(x)=x^(-2)sec(2/x)? i got -2x^-3(secxtanx2/x)

hartnn (hartnn):

you'll need product rule here...

hartnn (hartnn):

also you forgot to differentiate 2/x ....